# Predicting Critical Heat Flux with Uncertainty Quantification and Domain Generalization Using Conditional Variational Autoencoders and Deep Neural Networks

### v1

@@ 5.3 Results of Domain Generalization Analysis
In this subsection, we will evaluate the performance of the CVAE generative model and DNN predictive model under conditions outside the training data domain. This was done by comparing the errors in their outputs based on conditions from both inside and outside the training data domain. The testing set used in this work consist of 2,458 data points, with 1,047 points falling within the training data domain and the remaining 1,411 points falling outside the domain, as determined by the convex hull defined by the training dataset. The domain generalization analysis was performed using the mean CHF values obtained from the ensemble of DNNs and, for the CVAE, the mean value from generating multiple outputs. The relative error between $\mu_{\text{samples}}$ and the true value was calculated for the points falling inside and outside the training domain. The distributions of the relative errors for these two subsets and two models are shown in Figure 6.
@@
As shown in Figure 6, the error distributions for the two subsets (inside and outside the convex hull) behave similarly, with small deviations observed in the mean and standard deviation values. The maximum error values occur when predictions were made under conditions outside the training data domain. This applies to both the CVAE and DNN models. The mean, maximum and standard deviation values were calculated for the absolute relative errors for both models and subsets. These values, along with the fraction of points with error values greater than 10% in each subset are listed in Table 3.
@@
$\mu_{\text{error}}$ 1.2295 % 1.6653 %
$\text{Max}_{\text{error}}$ 9.8421 % 22.334 %
$\text{Std}_{\text{error}}$ 1.1734 % 2.2038 %
$F_{\text{error}}>10\%$ 0.0 % 0.9922 %
@@
$\mu_{\text{error}}$ 0.7364 % 0.9983 %
$\text{Max}_{\text{error}}$ 7.143 % 37.31 %
$\text{Std}_{\text{error}}$ 0.8139 % 1.8407 %
$F_{\text{error}}>10\%$ 0.0 % 0.6442 %
Comparing the mean error inside and outside the training domain, one can observe that there is an increase in the mean relative error for both models. However, the increase is not significant, with mean error values remaining less than 1.7% for both models and both subsets. A similar trend is observed for the standard deviation and the maximum of the relative errors. This behaviour is expected since extrapolation is being performed to unseen data domains. Notably, both models demonstrate favorable performance in predicting inside and outside the training domain. When predicting inside the training domain, none of the predictions resulted in errors greater than 10%, and only a minimal fraction of points have errors greater than 10% when predicting outside the training domain. This indicates that both models perform well in extrapolating to new domains, maintaining acceptable error values with mean errors smaller than 1.7%.
@@ 6 Discussions
The difference between the CVAE generative model and the DNN predictive model lies in their distinct approaches: DNN learns the input-output mapping of the training data in a supervised way, while CVAE learns the underlying distribution of the training data in an unsupervised way and generates synthetic data by sampling from the learned distribution. Since CVAE does not learn direct mapping between the input and output, one would expect it to perform less accurate when compared to a fine-tuned DNN model that is trained for regression tasks. However, our findings showed that the CVAE generative model can be powerful in generating synthetic CHF values, even when compared to DNN. The CVAE model maintained favorable values in all error metrics when compared to the results of a single DNN model.
@@
The UQ analysis of the CVAE model showed that its generative predictions have small variations at a specific input. This is demonstrated by the very small mean relative standard deviations of its predictions, which is 0.2579%. Additionally, changes in all metric values were very small when comparing the UQ and non-UQ resuts, as shown in Tables 1 and 2. These indicate that the CVAE model has low sensitivity to the random samples from the latent vector during the generation process. The CVAE model is thus robust, consistently maintaining similar error behavior across multiple predictions. This consistency indicates that, although the model produces slightly different outputs each time, the predictions will remain within the quantified error ranges, ensuring reliable performance even with variations in the latent vectors.
@@
On the other hand, quantifying uncertainties for the DNN model requires training multiple DNNs and measuring the variations between their predictions, which is more laborious than the CVAE model. The ensembled results showed significant improvement in all error metrics, except for the maximum error value, as shown in Tables 1 and 2.
This improvement can be attributed to the DNN model’s high sensitivity to weight initialization during training. Averaging over an ensemble of DNNs can thus produce more accurate predictions.
A ML model is generally expected to perform well for interpolation tasks, that is, when the model is applied to cases within the training domain. The ability of a ML model to extrapolate to new domains is important, especially if we aim to expand an existing dataset or test new experimental conditions.
@@
The domain generalization analysis showed that the CVAE and DNN models are reliable in generating and predicting CHF values outside the training domain. Both models showed consistent results when predicting both inside and outside the training domain. The mean relative absolute errors were smaller when predicting within the training domain, with a slight increase when predicting outside the training domain. The maximum relative absolute errors occurred when predicting outside the training domain. However, both models maintained their favorable performance in predicting CHF values within and outside the domain. With no predicted points having an absolute relative error greater than 10% within the training domain for both models, and less than 1% of points having an error greater than 10% when predicting outside the training domain. This indicates that both models were successful in extrapolating to new, unseen domains.
@@ 7 Conclusions
This study investigated the use of a Conditional Variational Autoencoders (CVAE) model in generation of critical heat flux (CHF) values and compared its performance with a traditional deep neural network (DNN) predictive model. This work was performed using the publicly-available CHF dataset used to construct the 2006 Groeneveld lookup table, consisting of nearly $25{,}000$ data points. Both the CVAE and DNN models were trained and tested using identical data partitions to ensure a direct comparison between them. The CVAE generative model demonstrated favorable test results across all error metrics when compared to a standalone DNN predictive model.
Uncertainty quantification (UQ) was then performed for both models to evaluate their robustness and consistency. For the CVAE model, we generated 200 samples from the latent vector and combined them with each input entry in the testing dataset before entering the decoder to generate a CHF value. In this way, the CVAE model produces 200 CHF values for each data point in the testing dataset, which can be used to get the uncertainty in the CVAE generated data and provide insight into the variability of the model’s predictions. For the DNN model, an ensemble was created by training 20 models of identical architecture with different initializations of the weights and biases. Similar to the CVAE, this produces a set of 20 “samples” for each data point in the testing dataset. The means and relative standard deviations were then taken along the samples of each testing data point. The UQ analysis revealed that the CVAE model had significantly smaller relative standard deviations when compared to the DNN ensemble. The CVAE generative model thus has a higher confidence and lower variability in its outputs than the DNN predictive model.
In the subsequent domain generalization analysis, the sampled CVAE and DNN ensemble’s outputs were then segmented into those inside and outside of the training set’s domain to evaluate how their performance changes when extrapolating. Both CVAE and DNN models exhibited strong extrapolation capabilities to the data outside the training domain. Inside the training domain, the mean error for the CVAE was $1.23$%, and for DNN ensemble, it was $0.74$%. Outside the training domain, the mean error increased to $1.67$% for the CVAE model and $1.00$% for the DNN ensemble. Both models performed generally well, with no predictions exceeding a $10$% absolute relative error within the training domain and less than $1$% of predictions exceeding this threshold outside the training domain.

### v2

@@ 5.3 Results of Domain Generalization Analysis
In this subsection, we will evaluate the performance of the CVAE generative model and DNN predictive model under conditions outside the training data domain. This was done by comparing the errors in their outputs based on conditions from both inside and outside the training data domain. The testing set used in this work consist of 2,458 data points, with 2,307 points falling within the training data domain and the remaining 151 points falling outside the domain, as determined by the convex hull defined by the training dataset. The domain generalization analysis was performed using the mean CHF values obtained from the ensemble of DNNs and, for the CVAE, the mean value from generating multiple outputs. The relative error between $\mu_{\text{samples}}$ and the true value was calculated for the points falling inside and outside the training domain. The distributions of the relative errors for these two subsets and two models are shown in Figure 6.
@@
[…]
@@
$\mu_{\text{error}}$ 7.1442 % 9.0927 %
$\text{Max}_{\text{error}}$ 111.39 % 67.192 %
$\text{Std}_{\text{error}}$ 9.114 % 11.399 %
$F_{\text{error}}>10\%$ 23.320 % 30.263 %
@@
$\mu_{\text{error}}$ 7.6951 % 11.625 %
$\text{Max}_{\text{error}}$ 118.501 % 116.695 %
$\text{Std}_{\text{error}}$ 9.8447 % 16.503 %
$F_{\text{error}}>10\%$ 25.498 % 34.868 %
Comparing the mean error inside and outside the training domain, one can observe that there is an increase in the mean relative error for both models. However, the increase is not significant, with mean error values remaining less than 12% for both models and both subsets. A similar trend is observed for the standard deviation of the relative errors. This behavior is expected since extrapolation is being performed to unseen data domains. Notably, both models demonstrate favorable performance in predicting inside and outside the training domain. When predicting inside the training domain, fewer predictions resulted in errors greater than 10%, while a slightly larger fraction of points exhibited errors above 10% when predicting outside the training domain. This suggests that both models generalize well to new domains, maintaining acceptable error levels with mean errors below 12%.
@@ 6 Discussions
The difference between the CVAE generative model and the DNN predictive model lies in their distinct approaches: DNN learns the input-output mapping of the training data in a supervised way, while CVAE learns the underlying distribution of the training data in an unsupervised way and generates synthetic data by sampling from the learned distribution. Since CVAE does not learn direct mapping between the input and output, one would expect it to perform less accurate when compared to a fine-tuned DNN model that is trained for regression tasks. However, our findings showed that the CVAE generative model can be powerful in generating synthetic CHF values, even when compared to DNN. The CVAE model maintained favorable values in all error metrics when compared to the DNN model.
@@
The UQ analysis of the CVAE model showed that its generative predictions have small variations at a specific input. This is demonstrated by the very small mean relative standard deviations of its predictions, which is 0.6187%. Additionally, changes in all metric values were very small when comparing the UQ and non-UQ resuts, as shown in Tables 1 and 2. These indicate that the CVAE model has low sensitivity to the random samples from the latent vector during the generation process. The CVAE model is thus robust, consistently maintaining similar error behavior across multiple predictions. This consistency indicates that, although the model produces slightly different outputs each time, the predictions will remain within the quantified error ranges, ensuring reliable performance even with variations in the latent vectors.
@@
On the other hand, quantifying uncertainties for the DNN model requires training multiple DNNs and measuring the variations between their predictions, which is more laborious than the CVAE model. The ensembled results showed improvement in all error metrics, except for the maximum error value, as shown in Tables 1 and 2. This improvement can be attributed to the DNN model’s high sensitivity to weight initialization during training. Averaging over an ensemble of DNNs can thus produce more accurate predictions. An ML model is generally expected to perform well for interpolation tasks, that is, when the model is applied to cases within the training domain. The ability of an ML model to extrapolate to new domains is important, especially if we aim to expand an existing dataset or test new experimental conditions.
@@
The domain generalization analysis showed that the CVAE and DNN models are reliable in generating and predicting CHF values outside the training domain. Both models showed consistent results when predicting both inside and outside the training domain. The mean relative absolute errors were smaller when predicting within the training domain, with a slight increase when predicting outside the training domain. However, both models maintained their favorable performance in predicting CHF values within and outside the domain. This indicates that both models were successful in extrapolating to new, unseen domains.
@@ 7 Conclusions
This study investigated the use of a Conditional Variational Autoencoders (CVAE) model in generation of critical heat flux (CHF) values and compared its performance with a traditional deep neural network (DNN) predictive model. This work was performed using the publicly-available CHF dataset used to construct the 2006 Groeneveld lookup table, consisting of nearly $25{,}000$ data points. Both the CVAE and DNN models were trained and tested using identical data partitions to ensure a direct comparison between them. The CVAE generative model demonstrated favorable test results across all error metrics when compared to the DNN predictive model.
Uncertainty quantification (UQ) was then performed for both models to evaluate their robustness and consistency. For the CVAE model, we generated 500 samples from the latent vector and combined them with each input entry in the testing dataset before entering the decoder to generate a CHF value. In this way, the CVAE model produces 500 CHF values for each data point in the testing dataset, which can be used to get the uncertainty in the CVAE generated data and provide insight into the variability of the model’s predictions. For the DNN model, an ensemble was created by training 20 models of identical architecture with different initializations of the weights and biases. Similar to the CVAE, this produces a set of 20 “samples” for each data point in the testing dataset. The means and relative standard deviations were then taken along the samples of each testing data point. The UQ analysis revealed that the CVAE model had significantly smaller relative standard deviations when compared to the DNN ensemble. The CVAE generative model thus has a higher confidence and lower variability in its outputs than the DNN predictive model.
In the subsequent domain generalization analysis, the sampled CVAE and DNN ensemble’s outputs were then segmented into those inside and outside of the training set’s domain to evaluate how their performance changes when extrapolating. Both CVAE and DNN models exhibited strong extrapolation capabilities to the data outside the training domain. Inside the training domain, the mean error for the CVAE was 7.14%, and for DNN ensemble, it was 7.69%. Outside the training domain, the mean error increased to 9.09% for the CVAE model and 11.62% for the DNN ensemble. Both models performed generally well when predicting/generating CHF both inside and outside the training domain.